A marketing research firm wishes to study the relationship between wine consumption and whether a person likes to watch professional tennis on television. One hundred randomly selected people are asked whether they drink wine and whether they watch tennis. The following results are obtained: Do Not Watch Tennis Watch TennisTotals Drink Wine Do Not Drink Wine Totals 28 52 80 37 63 100 20 (a) For each row and column total, calculate the coresponding row or column percentage Row 1 Row 2 Column 1 Columin2 (b) For each cell, calculate the corresponding cell, row, and column percentages. (Round your answers to the nearest whole number.) Watch Tennis Do Not Watch Tennis Drink Wine Cell- Do Not Drink Wine Cell- Cell- (c) Test the hypothesis that whether people drink wine is independent of whether people watch tennis. Set a .05. (Round your answer to 3 decimal places.) (Click to select. H). Conclude that whether a person drinks wine and whether a person watches tennis are | {Click to select) , events.
Answer #1

(a)

Row1 percentage:

(37 /100)*100% = 37%

Row2 percentage:

(63 /100)*100% = 63%

Column1 percentage:

(20 /100)*100% = 20%

Column2 percentage:

(80 /100)*100% = 80%

(b)

Following is the completed table:

Watch Tennis   Do not watch Tennis  
Drink Wine Observed   9 28
col (9/20)=45% (28/80)=35%
row 9/37=24% (28/37)=76%
column (9/100)=9% (28/100)=28%
Do not Drink wine Observed   11 52
col 55.0% 65.0%
row 17.0% 83.0%
column 11.0% 52.0%

(c)

Following is the output of chi square test generated by excel:

Chi-square Contingency Table Test for Independence Watch tennis Do not watch tennis 28 29.60 52 50.40 80 80.00 Total 37 37.00 63 63.00 100 100.00 Drink wine Observed Expected Do not drink wineObserved Expected Observed Expected 7.40 12.60 20 20.00 Total 69 chi-square 1 df 4074 p-value

The p-value is 0.4074

Since p-value is not less than a-0.05 so we fail to reject the null hypothesis.

That is we can conclude that events person drinks and person watches tennis are independent events.